Accelerator Update for April 29, 2009 - May 1, 2009 |
Wednesday April 29 The day shift began with Operations monitoring the Tevatron (TeV) store 7006 (98.0E30) and the Recycler stash (157.0E10), with the Antiproton Source stacking (16.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Meson MTest, MiniBooNE, and NuMI experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams. MI personnel conducted studies towards the end of the day shift. Operations terminated store 7006 at 5:45 PM and put the TeV into HEP shot setup. Operations established store 7007 at 7 PM with an initial luminosity of 319.0E30. At 8:10 PM, Operations reported the loss of two antiproton transfers due to an errant setting to a quadrupole power supply (M:Q103PB) from the Pbar sequencer. Pbar personnel investigated. Pbar, NuMI, MiniBooNE, and MTest took beam when available. Pbar transfers to the Recycler were completed when appropriate. Thursday April 30 The midnight shift began with Operations monitoring store 7007 (150.4E30) and the stash (161.6E10), and with Pbar stacking (16.6E10). Pbar sequencer problems continued during antiproton transfers with multiple settings and control issues for the Accumulator injection shutter. Operations terminated store 7007 at 10:30 AM and put the TeV into HEP shot setup. Operations established store 7008 at 12:30 PM with an initial luminosity of 327.0E30. Linac personnel tuned on a RF station (LRF4). The intensity of the I- Source decreased, which prompted a warm up of the H- Source in anticipation of a source switch. Booster intensities were adjusted, but the switch was never made. During the evening shift, Linac and Booster personnel tuned up for increased intensities. Pbar, NuMI, MiniBooNE, and MTest took beam when available. Pbar transfers to the Recycler were completed when appropriate. Friday May 1 The midnight shift began with Operations monitoring store 7008 (93.36E30) and the stash (298.7E10), and with Pbar stacking (12.3E10). The output intensity of the I- source decreased shortly after midnight. Operations contacted Linac personnel. Efforts to recover source intensities continued throughout the midnight shift. Operations terminated store 7008 at 4:15 AM and put the TeV into HEP shot setup. Operations established store 7009 at 5:40 AM with an initial luminosity of 332.0E30. Pbar, NuMI, MiniBooNE, and MTest took beam when available. Pbar transfers to the Recycler were completed when appropriate. The Plans for Friday and the Weekend The plans for today and the weekend are to stack, stash, and store. However, Operations will terminate store 7009 at 4 PM and the Run Coordinator will allow a four-hour access into the TeV sectors A-E for vacuum and separator work. |
|
More Information Comments and Suggestions |
| last modified 05/01/2009 email Fermilab |
| Security, Privacy, Legal | |